OpenStudy (anonymous):

Translate to a system of equations and solve. Two cars leave town at the same time going in opposite directions.One car travels 50 mi/h and the other travels 55 mi/h.In how many hours will they be 420 miles apart?

OpenStudy (anonymous):

55h + 50h = 420 105h = 420 h = 4

OpenStudy (anonymous):

let in t hours it will happen . in t hours 1st car will go 50t and second will go 55t. so total distance=(55+50)t hence 105t=420 t=4hrs